OpenStudy (anonymous):

-11p-4(4-5p)=5(p-3)-21

OpenStudy (anonymous):

i know i must distribute so does it equal -11p-16+20p=5p-15-2?

OpenStudy (anonymous):

21 at the end i mean

OpenStudy (nikato):

yes. that is correct. next, u want to combine like terms

OpenStudy (anonymous):

so 9p-16=5p-17?

OpenStudy (anonymous):

9p-16=5p=36 i mean

OpenStudy (nikato):

close. 5p-36

OpenStudy (anonymous):

this is the part that confuses me where do i go from there?

OpenStudy (nikato):

okay now subtract 5p on both isdes

OpenStudy (anonymous):

9p=5p-20?

OpenStudy (anonymous):

i solved it p = -5

OpenStudy (nikato):

ok so 4p=-20 and p=-5. yes. u r correct